### v1

# Risk Bounds for Mixture Density Estimation on Compact Domains via the $h$-Lifted Kullback–Leibler Divergence

Abstract
We consider the problem of estimating probability density functions based on sample data, using a finite mixture of densities from some component class. To this end, we introduce the $h$-lifted Kullback–Leibler (KL) divergence as a generalization of the standard KL divergence and a criterion for conducting risk minimization. Under a compact support assumption, we prove an $\mathcal{O}(1/{\sqrt{n}})$ bound on the expected estimation error when using the $h$-lifted KL divergence, which extends the results of Rakhlin et al. (2005, ESAIM: Probability and Statistics, Vol. 9) and Li and Barron (1999, Advances in Neural Information Processing Systems, Vol. 12) to permit the risk bounding of density functions that are not strictly positive. We develop a procedure for the computation of the corresponding maximum $h$-lifted likelihood estimators ($h$-MLLEs) using the Majorization-Maximization framework and provide experimental results in support of our theoretical bounds.
Keywords: Kullback–Leibler divergence, mixture density estimation, maximum likelihood, convergence rates, learning theory.
1 Introduction
Let $\left(\Omega,\mathfrak{A},{\mathrm{\mathbf{P}}}\right)$ be an abstract probability space and let $X:\Omega\to\mathcal{X}$ be a random variable taking values in the measurable space $\left(\mathcal{X},\mathfrak{F}\right)$, where $\mathcal{X}$ is a compact metric space equipped with its Borel $\sigma$-algebra $\mathfrak{F}$. Suppose that we observe an independent and identically distributed (i.i.d.) sample of random variables $\mathbf{X}_{n}=\left(X_{i}\right)_{i\in\left[n\right]}$, where $\left[n\right]=\left\{1,\ldots,n\right\}$, and that each $X_{i}$ arises from the same data generating process as $X$, characterized by the probability measure $F\ll\mu$ on $\left(\mathcal{X},\mathfrak{F}\right)$, with density function $f=\mathrm{d}F/\mathrm{d}\mu$, for some $\sigma$-finite $\mu$.
In this work, we are concerned with the estimating $f$ via a data dependent double-index sequence of estimators $\left(f_{k,n}\right)_{k,n\in\mathbb{N}}$, where
$f_{k,n}\in\mathcal{C}_{k}$ $=\mathrm{co}_{k}\left(\mathcal{P}\right)$
=$=\left\{f_{k}\left(\cdot;\psi_{k}\right)=\sum_{j=1}^{k}\pi_{j}\varphi\left(\cdot;\theta_{j}\right)\mid\varphi\left(\cdot;\theta_{j}\right)\in\mathcal{P},\,\pi_{j}\geq 0,\,j\in\left[k\right],\,\sum_{j=1}^{k}\pi_{j}=1\right\},$
for each $k,n\in\mathbb{N}$, and where
$\mathcal{P}=\left\{\varphi\left(\cdot;\theta\right):\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}\mid\theta\in\Theta\subset\mathbb{R}^{d}\right\},$
$\psi_{k}=\left(\pi_{1},\dots,\pi_{k},\theta_{1},\dots,\theta_{k}\right)$, and $d\in\mathbb{N}$. To ensure the measurability and existence of various optima, we shall assume that $\varphi$ is Caratheodory in the sense that $\varphi\left(\cdot;\theta\right)$ is $\left(\mathcal{X},\mathfrak{F}\right)$-measurable, for each $\theta\in\Theta$, and $\varphi\left(X;\cdot\right)$ is continuous for each $X\in\mathcal{X}$.
In the definition above, we can identify the set $\mathcal{C}_{k}=\mathrm{co}_{k}\left(\mathcal{P}\right)$ as the set of density functions that can be written as a convex combination of $k$ elements of $\mathcal{P}$, where $\mathcal{P}$ is often called the space of component density functions. We then interpret $\mathcal{C}_{k}$ as the class of $k$-component finite mixtures of densities of class $\mathcal{P}$, as studied, for example, by McLachlan and Peel (2004); Nguyen et al. (2020, 2022b).
1.1 Risk bounds for mixture density estimation
We are particularly interested in oracle bounds of the form
${\mathrm{\mathbf{E}}}\left\{\ell\left(f,f_{k,n}\right)\right\}-\ell\left(f,\mathcal{C}\right)\leq\rho\left(k,n\right),$ (1)
where $\left(p,q\right)\mapsto\ell\left(p,q\right)\in\mathbb{R}_{\geq 0}$ is a loss function on pairs of density functions. We also define the density-to-class loss
$\ell\left(f,\mathcal{C}\right)=\inf_{q\in\mathcal{C}}\ell\left(f,q\right),\ \mathcal{C}=\mathrm{cl}\left(\bigcup_{k\in\mathbb{N}}\mathrm{co}_{k}\left(\mathcal{P}\right)\right),$
where $\mathrm{cl}(\cdot)$ is the closure. Here, we identify $\left(k,n\right)\mapsto\rho\left(k,n\right)$ as a characterization of the rate at which the left-hand side of (1) converges to zero as $k$ and $n$ increase. Our present work follows the research of Li and Barron (1999), Rakhlin et al. (2005) and Klemelä (2007) (see also Klemelä 2009, Ch. 19). In Li and Barron (1999) and Rakhlin et al. (2005), the authors consider the case where $\ell\left(p,q\right)$ is taken to be the Kullback–Leibler (KL) divergence
$\mathrm{KL}\left(p\,||\,q\right)=\int p\log\frac{p}{q}\mathrm{d}\mu,$
and $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$ is a maximum likelihood estimator (MLE), where
${\psi}_{k,n}\in\underset{\psi_{k}\in\mathcal{S}_{k}\times\Theta^{k}}{\operatorname*{arg\,max\,}}\frac{1}{n}\sum_{i=1}^{n}\log f_{k}\left(X_{i};\psi_{k}\right),$
is a function of $\mathbf{X}_{n}$, with $\mathcal{S}_{k}$ denoting the probability simplex in $\mathbb{R}^{k}$.
Under the assumption that $f,f_{k}\geq a$, for some $a>0$ and each $k\in\left[n\right]$ (i.e., strict positivity), Li and Barron (1999) obtained the bound
${\mathrm{\mathbf{E}}}\left\{\mathrm{KL}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}\left(f\,||\,\mathcal{C}\right)\leq c_{1}\frac{1}{k}+c_{2}\frac{k\log\left(c_{3}n\right)}{n},$
for constants $c_{1},c_{2},c_{3}>0$, which was then improved by Rakhlin et al. (2005) who obtain the bound
${\mathrm{\mathbf{E}}}\left\{\mathrm{KL}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}\left(f\,||\,\mathcal{C}\right)\leq c_{1}\frac{1}{k}+c_{2}\frac{1}{\sqrt{n}},$
for constants $c_{1},c_{2}>0$ (constants $(c_{j})_{j\in\mathbb{N}}$ are typically different between expressions).
Alternatively, Klemelä (2007) takes $\ell\left(p,q\right)$ to be the squared $L_{2}\left(\mu\right)$ norm distance (i.e., the least-squares loss):
$\ell\left(p,q\right)=\lVert p-q\rVert_{2,\mu}^{2},$
where $\lVert p\rVert_{2,\mu}^{2}=\int_{\mathcal{X}}\lvert p\rvert^{2}\mathrm{d}\mu$, for each $p\in L_{2}\left(\mu\right)$, and choose $f_{k,n}$ as minimizers of the $L_{2}\left(\mu\right)$ empirical risk: i.e., $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
$\psi_{k,n}\in\underset{\psi_{k}\in\mathcal{S}_{k}\times\Theta^{k}}{\operatorname*{arg\,min\,}}-\frac{2}{n}\sum_{i=1}^{n}f_{k}\left(\cdot;\psi_{k}\right)+\left\|f_{k}\left(\cdot;\psi_{k}\right)\right\|_{2,\mu}^{2}.$ (2)
Here, Klemelä (2007) establish the bound
${\mathrm{\mathbf{E}}}\left\|f-{f}_{k,n}\right\|_{2,\mu}^{2}-\inf_{q\in\mathcal{C}}\left\|f-q\right\|_{2,\mu}^{2}\leq c_{1}\frac{1}{k}+c_{2}\frac{1}{\sqrt{n}},$
$c_{1},c_{2}>0$, without the lower bound assumption on $f,f_{k}$ above, even permitting $\mathcal{X}$ to be unbounded. Via the main results of Naito and Eguchi (2013), the bound above can be generalized to the $U$-divergences, which includes the special $L_{2}(\mu)$ norm distance as a special case.
On the one hand, the sequence of MLEs required for the results of Li and Barron (1999) and Rakhlin et al. (2005) are typically computable, for example, via the usual expectation–maximization approach (cf. McLachlan and Peel 2004, Ch. 2). This contrasts with the computation of least-squares density estimators of form (2), which requires evaluations of the typically intractable integral expressions: $\left\|f_{k}\left(\cdot;\psi_{k}\right)\right\|_{2}^{2}$. However, the least-squares approach of Klemelä (2007) permits the analysis using families $\mathcal{P}$ of usual interest, such as normal distributions and beta distributions, the latter of which being compactly supported but having densities that cannot be bounded away from zero without restrictions, and thus do not satisfy the regularity conditions of Li and Barron (1999) and Rakhlin et al. (2005).
1.2 Main contributions
We propose the following $h$-lifted KL divergence, as a generalization of the standard KL divergence to address the computationally tractable estimation of density functions which do not satisfy the regularity conditions of Li and Barron (1999) and Rakhlin et al. (2005). The use of the $h$-lifted KL divergence has the possibility to advance theories based on the standard KL divergence in statistical machine learning. To this end, let $h:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ be a function in $L_{1}(\mu)$, and define the $h$-lifted KL divergence by:
$\mathrm{KL}_{h}\left(p\,||\,q\right)=\int_{\mathcal{X}}\left\{p+h\right\}\log\frac{p+h}{q+h}\mathrm{d}\mu.$ (3)
In the sequel, we shall show that $\mathrm{KL}_{h}$ is a Bregman divergence on the space of probability density functions, as per Csiszár (1995).
Assume that $h$ is a probability density function, and let $\mathbf{Y}_{n}=\left(Y_{i}\right)_{i\in\left[n\right]}$ be a an i.i.d. sample, independent of $\mathbf{X}_{n}$, where each $Y_{i}:\Omega\rightarrow\mathcal{X}$ is a random variable with probability measure on $(\mathcal{X},\mathfrak{F})$, characterized by the density $h$ with respect to $\mu$. Then, for each $k$ and $n$, let $f_{k,n}$ be defined via the maximum $h$-lifted likelihood estimator ($h$-MLLE): $f_{k,n}=f_{k}\left(\cdot;{\psi}_{k,n}\right)$, where
${\psi}_{k,n}\in\underset{\psi_{k}\in\mathcal{S}_{k}\times\Theta^{k}}{\operatorname*{arg\,max\,}}~{}\frac{1}{n}\sum_{i=1}^{n}\left(\log\left\{f_{k}\left(X_{i};\psi_{k}\right)+h\left(X_{i}\right)\right\}+\log\left\{f_{k}\left(Y_{i};\psi_{k}\right)+h\left(Y_{i}\right)\right\}\right).$ (4)
The primary aim of this work is to show that
${\mathrm{\mathbf{E}}}\left\{\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)\leq c_{1}\frac{1}{k}+c_{2}\frac{1}{\sqrt{n}}$ (5)
for some constants $c_{1},c_{2}>0$, without requiring the strict positivity assumption that $f,f_{k}\geq a>0$.
This result is a compromise between the works of Li and Barron (1999) and Rakhlin et al. (2005), and Klemelä (2007), as it applies to a broader space of component densities $\mathcal{P}$, and because the required $h$-MLLEs, (4), can be efficiently computed via minorization–maximization (MM) algorithms (see, e.g., Lange 2016). We shall discuss this assertion in Section 6.
1.3 Relevant literature
Our work largely follows the approach of Li and Barron (1999), which have was extended upon by Rakhlin et al. (2005) and Klemelä (2007). All three texts use approaches based on the availability of greedy algorithms for maximizing convex functions with convex functional domains. In this work, we shall make use of the proof techniques of Zhang (2003). Related results in this direction can be found in DeVore and Temlyakov (2016) and Temlyakov (2016). Making the same boundedness assumption as Rakhlin et al. (2005), Dalalyan and Sebbar (2018) obtain refined oracle inequalities under the additional assumption that the class $\mathcal{P}$ is finite. Numerical implementations of greedy algorithms for estimating finite mixtures of Gaussian densities were studied by Vlassis and Likas (2002) and Verbeek et al. (2003).
The $h$-MLLE as an optimization objective can be compared to other similar modified likelihood estimators, such as the $L_{q}$ likelihood of Ferrari and Yang (2010) and Qin and Priebe (2013), the $\beta$-likelihood of Basu et al. (1998) and Fujisawa and Eguchi (2006), penalized likelihood estimators, such as maximum a posteriori estimators of Bayesian models, or $f$-separable Bregman distortion measures of Kobayashi and Watanabe (2024, 2021).
The practical computation of the $h$-MLLEs, (4), is made possible via the MM algorithm framework of Lange (2016), see also Hunter and Lange (2004), Wu and Lange (2010), and Nguyen (2017) for further details. Such algorithms have well-studied global convergence properties and can be modified for mini-batch and stochastic settings (see, e.g., Razaviyayn et al., 2013 and Nguyen et al., 2022a).
A related and popular setting of investigations is that of model selection, where the objects of interest are single-index sequences $\left(f_{k_{n},n}\right)_{n\in\mathbb{N}}$, and where the aim is to obtain finite-sample bounds for losses of the form $\ell\left(f_{k_{n},n},f\right)$, where each $k_{n}\in\mathbb{N}$ is a data dependent function, often obtained by optimizing some penalized loss criterion, as described in Massart (2007), Koltchinskii (2011, Ch. 6), and Giraud (2021, Ch. 2). In the context of finite mixtures, examples of such analyses can be found in the works of Maugis and Michel (2011) and Maugis-Rabusseau and Michel (2013). A comprehensive bibliography of model selection results for finite mixtures and related statistical models can be found in Nguyen et al. (2022c).
1.4 Organization of paper
The manuscript proceeds as follows. In the following section we formally characterize the $h$-lifted KL divergence as a Bregman divergence, and establish some of its properties. In Section 4, we prove new risk bounds of the form (1) in terms of the $h$-lifted KL divergence. We discuss the computation of the $h$-lifted likelihood estimator of form (4) in Section 6, followed by the presentation of some empirical results regarding the convergence of (1) in terms of both $k$ and $n$. Additional technical results are also included in the Appendices.
2 The $h$-lifted KL divergence and its properties
In this section we formally define the $h$-lifted KL divergence on the space of density functions and establish some of its properties.
Definition 1 ($h$-lifted KL divergence).
Let $f,g$, and $h$ be probability density functions on the space $\mathcal{X}$, where $h>0$. The $h$-lifted KL divergence from $g$ to $f$ is defined as follows:
$\mathrm{KL}_{h}\left(f\,||\,g\right)=\int_{\mathcal{X}}\left\{f+h\right\}\log\frac{f+h}{g+h}\mathrm{d}\mu={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\log\frac{f+h}{g+h}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\log\frac{f+h}{g+h}\right\}.$
2.1 $\textrm{KL}_{h}$ as a Bregman divergence
Let $\phi:\mathcal{I}\to\operatorname*{\mathbb{R}}$, $\mathcal{I}=(0,\infty)$ be a strictly convex function that is continuously differentiable. The Bregman divergence between scalars $d_{\phi}:\mathcal{I}\times\mathcal{I}\to\mathbb{R}_{\geq 0}$ generated by the function $\phi$ is given by:
$d_{\phi}(p,q)=\phi(p)-\phi(q)-\phi^{\prime}(q)(p-q),$
where $\phi^{\prime}(q)$ denotes the derivative of $\phi$ at $q$.
Bregman divergences possess several useful properties, including the following list:
  1. 1.
    Non-negativity: $d_{\phi}(p,q)\geq 0$ for all $p,q\in\mathcal{I}$ with equality if and only if $p=q$;
  2. 2.
    Asymmetry: $d_{\phi}(p,q)\neq d_{\phi}(q,p)$ in general;
  3. 3.
    Convexity: $d_{\phi}(p,q)$ is convex in $p$ for every fixed $q\in\mathcal{I}$.
  4. 4.
    Linearity: $d_{c_{1}\phi_{1}+c_{2}\phi_{2}}(p,q)=c_{1}\,d_{\phi_{1}}(p,q)+c_{2}\,d_{\phi_{2}}(p,q)$ for $c_{1},c_{2}\geq 0$.
The properties for Bregman divergences between scalars can be extended to density functions and other functional spaces, as established in Frigyik et al. (2008) and Stummer and Vajda (2012), for example. We also direct the interested reader to the works of Pardo (2006), Basu et al. (2011), and Amari (2016).
The class of $h$-lifted KL divergences constitute a generalization of the usual KL divergence and are a subset of the Bregman divergences over the space of density functions that are considered by Csiszár (1995). Namely, let $\mathcal{P}$ be a convex set of probability densities with respect to the measure $\mu$ on $\mathcal{X}$. The Bregman divergence $D_{\phi}:\mathcal{P}\times\mathcal{P}\to[0,\infty)$ between densities $p,q\in\mathcal{P}$ can be constructed as follows:
$D_{\phi}(p\,||\,q)=\int_{\mathcal{X}}d_{\phi}\left(p(x),q(x)\right)\mathrm{d}\mu(x).$
The $h$-lifted KL divergence $\mathrm{KL}_{h}$ as a Bregman divergence is generated by the function $\phi(u)=(u+h)\log(u+h)-(u+h)+1$. This assertion is demonstrated in Appendix A.
2.2 Advantages of the $h$-lifted KL divergence
When the standard KL divergence is employed in the density estimation problem, it is common to restrict consideration of density functions to those bounded away from zero by some positive constant. That is, one typically considers the smaller class of so-called admissible target densities $\mathcal{P}_{\alpha}\subset\mathcal{P}$ (cf. Meir and Zeevi, 1997), where
$\mathcal{P}_{\alpha}=\left\{\varphi(\cdot;\theta)\in\mathcal{P}\mid\varphi(\cdot;\theta)\geq\alpha>0\right\}.$
Without this restriction, the standard KL divergence can be unbounded, even for functions with bounded $L_{1}$ norms. For example, let $p$ and $q$ be densities of beta distributions on the support $\mathcal{X}=\left[0,1\right]$. That is, suppose that $p,q\in\mathcal{P}_{\mathrm{beta}}$, respectively characterized by parameters $\theta_{p}=\left(a_{p},b_{p}\right)$ and $\theta_{q}=\left(a_{q},b_{q}\right)$, where
$\mathcal{P}_{\mathrm{beta}}=\left\{x\mapsto\beta\left(x;\theta\right)=\frac{\Gamma\left(a+b\right)}{\Gamma\left(a\right)\Gamma\left(b\right)}x^{a-1}\left(1-x\right)^{b-1},\theta=\left(a,b\right)\in\mathbb{R}_{>0}^{2}\right\}.$ (6)
Then, from Gil et al. (2013), the KL divergence between $p$ and $q$ is given by:
$\mathrm{KL}\left(p\,||\,q\right)=\log\left\{\frac{\Gamma\left(a_{q}\right)\Gamma\left(b_{q}\right)}{\Gamma\left(a_{q}+b_{q}\right)}\right\}-\log\left\{\frac{\Gamma\left(a_{p}\right)\Gamma\left(b_{p}\right)}{\Gamma\left(a_{p}+b_{p}\right)}\right\}\\ +\left(a_{p}-a_{q}\right)\left\{\psi\left(a_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\}+\left(b_{p}-b_{q}\right)\left\{\psi\left(b_{p}\right)-\psi\left(a_{p}+b_{p}\right)\right\},$
where $\psi:\mathbb{R}_{>0}\rightarrow\mathbb{R}$ is the digamma function. Next, suppose that $a_{p}=b_{q}$ and $a_{q}=b_{p}=1$, which leads to the simplification
$\mathrm{KL}\left(p\,||\,q\right)=\left(a_{p}-1\right)\left\{\psi\left(a_{p}\right)-\psi(1)\right\}.$
Since $\psi$ is strictly increasing, we observe that the right-hand side diverges as $a_{p}\to\infty$. Thus, the KL divergence between beta distributions is unbounded. The $h$-lifted KL divergence in contrast does not suffer from this problem, and does not require the restriction to $\mathcal{P}_{\alpha}$. This allows us to consider cases where $p,q\in\mathcal{P}$ are not bounded away from 00.
Proposition 2.
$\mathrm{KL}_{h}\left(p\,||\,q\right)$ is bounded for all continuous densities $p,q\in\mathcal{P}$.
Proof.  Let $\tilde{p}=p+h$ and $\tilde{q}=q+h$. Since $h$ is positive, there exists some $\tilde{q}_{*}$ such that $\tilde{q}_{*}=\inf_{x\in\mathcal{X}}\left\{q(x)+h(x)\right\}>0$. Similarly, since $\mathcal{X}$ is compact, there exists some positive $\tilde{p}^{*}$ such that $0<\tilde{p}^{*}=\sup_{x\in\mathcal{X}}\left\{p(x)+h(x)\right\}<\infty$. Define $M=\sup_{x\in\mathcal{X}}\log\{\tilde{p}(x)/\tilde{q}(x)\}$. Then $M<\infty$, and
$\mathrm{KL}_{h}\left(p\,||\,q\right)=\int_{\mathcal{X}}\tilde{p}\log\frac{\tilde{p}}{\tilde{q}}\mathrm{d}\mu\leq\sup_{x\in\mathcal{X}}\log\frac{\tilde{p}}{\tilde{q}}\int_{\mathcal{X}}\tilde{p}\mathrm{d}\mu=2M<\infty.$
Let $L_{p}(f,g)$ denote the standard $L_{p}$ norm,
$L_{p}(f,g)=\left\{\int_{\mathcal{X}}\left\lvert f(x)-g(x)\right\rvert^{p}\mathrm{d}\mu(x)\right\}^{1/p}.$
As remarked previously, Klemelä (2007) established empirical risk bounds in terms of the $L_{2}$ norm distance. Following results from Meir and Zeevi (1997), characterizing the relationship between the KL divergence in terms of the $L_{2}$ norm distance, we establish the corresponding relationship between the $h$-lifted KL divergence and the $L_{2}$ norm distance.
Proposition 3.
For probability density functions $f,g$, and $h$, where $h$ is such that $h(x)\geq\gamma>0$ for all $x\in\mathcal{X}$, the following inequality holds:
$\mathrm{KL}_{h}\left(f\,||\,g\right)\leq\gamma^{-1}L_{2}^{2}(f,g).$
Proof.  Defining $\tilde{f}$ and $\tilde{g}$ as above, we have
$\mathrm{KL}_{h}\left(f\,||\,g\right)=\int_{\mathcal{X}}\tilde{f}\log\frac{\tilde{f}}{\tilde{g}}\mathrm{d}\mu\leq\int_{\mathcal{X}}\tilde{f}\left(\frac{\tilde{f}}{\tilde{g}}-1\right)\mathrm{d}\mu=\int_{\mathcal{X}}\frac{(f-g)^{2}}{\tilde{g}}\mathrm{d}\mu\leq\gamma^{-1}L_{2}^{2}(f,g),$
The following section is devoted to establishing some technical definitions and instrumental results which will be utilized in later sections.
3 Preliminaries
Recall that we are interested in bounds of the form (1). Note that $\mathcal{P}$ is a subset of the linear space
$\mathcal{V}=\text{cl}\left(\bigcup_{k\in\mathbb{N}}\left\{\sum_{j=1}^{k}\varpi_{j}\varphi\left(\cdot;\theta_{j}\right)\mid\varphi\left(\cdot;\theta_{j}\right)\in\mathcal{P},\varpi_{j}\in\mathbb{R},j\in\left[k\right]\right\}\right),$
and hence we can apply the following result, paraphrased from Zhang (2003, Thm. II.1).
Lemma 4.
Let $\kappa$ be a differentiable and convex function on $\mathcal{V}$, and let $\left(\bar{f}_{k}\right)_{k\in\mathbb{N}}$ be a sequence of functions obtained by Algorithm 1. If
$\sup_{p,q\in\mathcal{C},\pi\in\left(0,1\right)}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq\mathfrak{M}<\infty,$
then, for each $k\in\mathbb{N}$,
$\kappa\left(\bar{f}_{k}\right)-\inf_{p\in\mathcal{C}}\kappa\left(p\right)\leq\frac{2\mathfrak{M}}{k+2}.$
Algorithm 1 Algorithm for computing a greedy approximation sequence.
1:  Input: $\bar{f}_{0}\in\mathcal{P}$
2:  for $k\in\mathbb{N}$ do
3:     Compute $\left(\bar{\pi}_{k},\bar{\theta}_{k}\right)=\operatorname*{arg\,min\,}\limits_{\left(\pi,\theta\right)\in[0,1]\times\Theta}\kappa\left(\left(1-\pi\right)\bar{f}_{k-1}+\pi\varphi\left(\cdot;\theta\right)\right)$
4:     Define $\bar{f}_{k}=\left(1-\bar{\pi}_{k}\right)\bar{f}_{k-1}+\bar{\pi}_{k}\varphi\left(\cdot;\bar{\theta}_{k}\right)$
5:  end for
We are interested in two choices for $\kappa$:
$\kappa\left(p\right)=\mathrm{KL}_{h}\left(f\,||\,p\right)$ (7)
and its sample counterpart,
$\kappa_{n}\left(p\right)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{f\left(X_{i}\right)+h\left(X_{i}\right)}{p\left(X_{i}\right)+h\left(X_{i}\right)}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{f\left(Y_{i}\right)+h\left(Y_{i}\right)}{p\left(Y_{i}\right)+h\left(Y_{i}\right)},$ (8)
where $(X_{i})_{i\in[n]}$ and $(Y_{i})_{i\in[n]}$ are realisations of $X$ and $Y$, respectively. For choice (7), by the dominated convergence theorem, we observe that
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)$
$\qquad\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}$
$\qquad\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}.$
Suppose that each $\varphi\left(\cdot;\theta\right)\in\mathcal{P}$ is bounded from above by $c<\infty$. Then, since $p,q\in\mathcal{C}$ are non-negative functions, we have the fact that $\left(p-q\right)^{2}\leq c^{2}$. If we further have $a\leq h$ for some $a>0$, then $\left[\left(1-\pi\right)p+\pi q+h\right]^{2}\geq a^{2}$, which implies that
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq 2\times\frac{c^{2}}{a^{2}}$
for every $p,q\in\mathcal{C}$ and $\pi\in\left(0,1\right)$, and thus
$\sup_{p,q\in\mathcal{C},\pi\in\left(0,1\right)}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq\frac{2c^{2}}{a^{2}}<\infty.$
Similarly, for case (8), we have
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa_{n}\left(\left(1-\pi\right)p+\pi q\right)=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f\left(x_{i}\right)+h\left(x_{i}\right)}{\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)}\\ +\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f\left(y_{i}\right)+h\left(y_{i}\right)}{\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)}\\ =\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(x_{i}\right)-q\left(x_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)\right]^{2}}+\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(y_{i}\right)-q\left(y_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)\right]^{2}}.$
By the same argument, as for $\kappa$, we have the fact that $\left(p\left(x\right)-q\left(x\right)\right)^{2}\leq c^{2}$, for every $p,q\in\mathcal{C}$ and every $x\in\mathcal{X}$, and furthermore $\left[\left(1-\pi\right)p\left(x\right)+\pi q\left(x\right)+h\left(x\right)\right]^{2}\geq a^{2}$, for any $\pi\in\left(0,1\right)$. Thus,
$\sup_{p,q\in\mathcal{C},\pi\in\left(0,1\right)}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq\frac{2c^{2}}{a^{2}}<\infty,$
as required. We therefore obtain the following result.
Proposition 5.
Let $\square$ denote either $\kappa$, the $\textrm{KL}_{h}$ divergence (7), or $\kappa_{n}$, the sample $\textrm{KL}_{h}$ divergence (8), and assume that $h\geq a$ and $\varphi\left(\cdot;\theta\right)\leq c$, for each $\theta\in\Theta$. Then,
$\square\left(\bar{f}_{k}\right)-\inf_{p\in\mathcal{C}}\square\left(p\right)\leq\frac{4a^{-2}c^{2}}{k+2},$
for each $k\in\mathbb{N}$, where $\left(\bar{f}_{k}\right)_{k\in\mathbb{N}}$ is obtained as per Algorithm 1.
Notice that sequences $\left(\bar{f}_{k}\right)_{k\in\mathbb{N}}$ obtained via Lemma 4 are greedy approximation sequences, and that $\bar{f}_{k}\in\mathcal{C}_{k}$, for each $k\in\mathbb{N}$. Let $\left(f_{k}\right)_{k\in\mathbb{N}}$ be the sequence of minimizers defined by
$f_{k}=\underset{\psi_{k}\in\mathcal{S}_{k}\times\Theta^{k}}{\operatorname*{arg\,min\,}}\mathrm{KL}_{h}\left(f\,||\,f_{k}\left(\cdot;\psi_{k}\right)\right),$ (9)
and let $\left(f_{k,n}\right)_{k\in\mathbb{N}}$ be the sequence of $h$-MLLEs, as per (4). Then, by definition, we have the fact that $\kappa\left(f_{k}\right)\leq\kappa\left(\bar{f}_{k}\right)$ and $\kappa\left(f_{k,n}\right)\leq\kappa\left(\bar{f}_{k}\right)$, for $\kappa$ set as (7) or (8), respectively. Thus, we have the following result.
Proposition 6.
For the $\textrm{KL}_{h}$ divergence (7), under the assumption that $h\geq a$ and $\varphi\left(\cdot;\theta\right)\leq c$, for each $\theta\in\Theta$, we have
$\kappa\left(f_{k}\right)-\inf_{p\in\mathcal{C}}\kappa\left(p\right)\leq\frac{4a^{-2}c^{2}}{k+2}$ (10)
for each $k\in\mathbb{N}$, where $\left(f_{k}\right)_{k\in\mathbb{N}}$ is the sequence of minimizers defined via (9). Furthermore, for the sample $\textrm{KL}_{h}$ divergence (8), under the same assumptions as above, we have
$\kappa_{n}\left(f_{k,n}\right)-\inf_{p\in\mathcal{C}}\kappa_{n}\left(p\right)\leq\frac{4a^{-2}c^{2}}{k+2},$ (11)
for each $k\in\mathbb{N}$, where $\left(f_{k,n}\right)_{k\in\mathbb{N}}$ are $h$-MLLEs defined via (4).
As is common in many statistical learning/uniform convergence results (e.g., Bartlett and Mendelson, 2002, Koltchinskii and Panchenko, 2004), we employ the use of Rademacher processes and associated bounds. Let $(\varepsilon_{i})_{i\in[n]}$ be i.i.d. Rademacher random variables, that is $\mathrm{\mathbf{P}}(\varepsilon_{i}=-1)=\mathrm{\mathbf{P}}(\varepsilon_{i}=1)=\nicefrac{{1}}{{2}}$, that are independent of $(X_{i})_{i\in[n]}$. The Rademacher process, indexed by a class of real measurable functions $\mathcal{S}$, is defined as the quantity
$R_{n}(s)=\frac{1}{n}\sum_{i=1}^{n}s(X_{i})\varepsilon_{i},$
for $s\in\mathcal{S}$. The Rademacher complexity of the class $\mathcal{S}$ is given by $\mathcal{R}_{n}(\mathcal{S})=\operatorname*{{\mathrm{\mathbf{E}}}}\sup_{s\in\mathcal{S}}\lvert R_{n}(s)\rvert$.
In the subsequent section, we make use of the following result regarding the supremum of convex functions:
Lemma 7 (Rockafellar, 1997, Thm. 32.2).
Let $\eta$ be a convex function on a linear space $\mathcal{T}$, and let $\mathcal{S}\subset\mathcal{T}$ be an arbitrary subset. Then,
$\sup_{p\in\mathcal{S}}\eta\left(p\right)=\sup_{p\in\mathrm{co}\left(\mathcal{S}\right)}\eta\left(p\right).$
In particular, we use the fact that since a linear functional of convex combinations achieves its maximum value at vertices, the Rademacher complexity of $\mathcal{S}$ is equal to the Rademacher complexity of $\textrm{co}(\mathcal{S})$ (see Lemma 19). We consequently obtain the following result.
Lemma 8.
Let $(\varepsilon_{i})_{i\in[n]}$ be i.i.d. Rademacher random variables, independent of $(X_{i})_{i\in[n]}$ and $\mathcal{P}$ be defined as above. The sets $\mathcal{C}$ and $\mathcal{P}$ will have equal complexity, $\mathcal{R}_{n}(\mathcal{C})=\mathcal{R}_{n}(\mathcal{P})$, and the supremum of the Rademacher process indexed by $\mathcal{C}$ is equal to the supremum on the basis functions of $\mathcal{P}$:
${\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\sup_{g\in\mathcal{C}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}g(X_{i})\varepsilon_{i}\right\rvert={\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\sup_{\theta\in\Theta}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varphi(X_{i};\theta)\varepsilon_{i}\right\rvert.$
Proof.  Follows immediately from Lemma 7.
4 Main results
Here we provide explicit statements regarding the convergence rates claimed in (5). We assume that $f$ is bounded above by some constant $c$ and that the lifting function $h$ is bounded above and below by constants $a$ and $b$, respectively.
Theorem 9.
Let $h$ be a positive density satisfying $0<a\leq h(x)\leq b$, for all $x\in\mathcal{X}$. For any target density $f$ satisfying $0\leq f(x)\leq c$, for all $x\in\mathcal{X}$ and where $f_{k,n}$ is the minimizer of $\mathrm{KL}_{h}$ over $k$-component mixtures, the following inequality holds:
$\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)\leq\frac{u_{1}}{k+2}+\frac{u_{2}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{u_{3}}{\sqrt{n}},$
where $u_{1}$, $u_{2}$, and $u_{3}$ are positive constants that depend on some or all of $a$, $b$, and $c$.
Corollary 10.
Let $\mathcal{X}$ and $\Theta$ be compact and assume the following Lipschitz condition holds: for each $x\in\mathcal{X}$, and for each $\theta,\tau\in\Theta$,
$\left|\varphi\left(x;\theta\right)-\varphi\left(x;\tau\right)\right|\leq\varPhi\left(x\right)\left\|\theta-\tau\right\|_{1},$ (12)
for some function $\varPhi:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$, where $\lVert\varPhi\rVert_{\infty}=\sup_{x\in\mathcal{X}}\lvert\varPhi(x)\rvert<\infty$. Then the bound in Theorem 9 becomes
$\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)\leq\frac{c_{1}}{k+2}+\frac{c_{2}}{\sqrt{n}},$
where $c_{1}$ and $c_{2}$ are positive constants.
5 Proofs
We first present a result establishing a uniform concentration bound for the $h$-lifted log-likelihood ratios, which is instrumental in the proof of Theorem 9. Our proofs broadly follow the structure of Rakhlin et al. (2005), modified as needed for the use of $\textrm{KL}_{h}$.
Assume that $0\leq\varphi(\cdot;\theta)<{c}$ for some $c\in\operatorname*{\mathbb{R}}_{>0}$. For brevity, we adopt the notation: $\left\lVert T(g)\right\rVert_{\mathcal{C}}=\sup_{g\in\mathcal{C}}\lvert T(g)\rvert$.
Theorem 11.
Let $X_{1},\ldots,X_{n}$ be an i.i.d. sample of size $n$ drawn from a fixed density $f$ such that $0\leq f(x)\leq c$ for all $x\in\mathcal{X}$, and let $h$ be a positive density with $0<a\leq h(x)\leq b$ for all $x\in\mathcal{X}$. Then with probability at least $1-\mathrm{e}^{-t}$,
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{g(X_{i})+h(X_{i})}{f(X_{i})+h(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{g+h}{f+h}\right\rVert_{\mathcal{C}}\leq\frac{w_{1}}{\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}},$
where $w_{1}$, $w_{2}$, and $w_{3}$ are constants that each depend on some or all of $a$, $b$, and $c$, and $N(\mathcal{P},\varepsilon,d_{n,x})$ is the $\varepsilon$-covering number of $\mathcal{P}$ with respect to the following empirical $L_{2}$ metric
$d_{n,x}^{2}(\varphi_{1},\varphi_{2})=\frac{1}{n}\sum_{i=1}^{n}(\varphi_{1}(X_{i})-\varphi_{2}(X_{i}))^{2}.$
Remark 12.
The bound on the term
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{g(Y_{i})+h(Y_{i})}{f(Y_{i})+h(Y_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{g+h}{f+h}\right\rVert_{\mathcal{C}}$
is the same as the above, except where the empirical distance $d_{n,x}$ is replaced by $d_{n,y}$, defined in the same way as $d_{n,x}$ but with $Y_{i}$ replacing $X_{i}$.
Proof. [of Theorem 11]. Fix $h$ and define the following quantities: $\tilde{g}=g+h$, $\tilde{f}=f+h$, $\tilde{\mathcal{C}}=\mathcal{C}+h$,
$m_{i}=\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})},\quad m_{i}^{\prime}=\log\frac{\tilde{g}(X_{i}^{\prime})}{\tilde{f}(X_{i}^{\prime})},\quad Z(x_{1},\ldots,x_{n})=\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-\operatorname*{{\mathrm{\mathbf{E}}}}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert_{\tilde{\mathcal{C}}}.$
We first apply McDiarmid’s inequality (Lemma 21) to the random variable $Z$. The bound on the martingale difference is given by
$\left\lvert Z(X_{1},\ldots,X_{i},\ldots,X_{n})-Z(X_{1},\ldots,X_{i}^{\prime},\ldots,X_{n})\right\rvert$ $=\left\lvert\left\lVert\operatorname*{{\mathrm{\mathbf{E}}}}\log\frac{\tilde{g}}{\tilde{f}}-\frac{1}{n}(m_{1}+\!...\!+m_{i}+\!...\!+m_{n})\right\rVert_{\tilde{\mathcal{C}}}\right.$
$-\left.\left\lVert\operatorname*{{\mathrm{\mathbf{E}}}}\log\frac{\tilde{g}}{\tilde{f}}-\frac{1}{n}(m_{1}+\!...\!+m_{i}^{\prime}+\!...\!+m_{n})\right\rVert_{\tilde{\mathcal{C}}}\right\rvert$
$\leq\frac{1}{n}\left\lVert\log\frac{\tilde{g}(X_{i}^{\prime})}{\tilde{f}(X_{i}^{\prime})}-\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\right\rVert_{\tilde{\mathcal{C}}}$
$\leq\frac{1}{n}\left(\log\frac{c+b}{a}-\log\frac{a}{c+b}\right)$
$=\frac{1}{n}2\log\frac{c+b}{a}=c_{i}.$
The chain of inequalities holds because of the triangle inequality and the properties of the supremum. By Lemma 21, we have
${\mathrm{\mathbf{P}}}(Z-\operatorname*{{\mathrm{\mathbf{E}}}}Z>\varepsilon)\leq\exp\left\{-\frac{n\varepsilon^{2}}{(\sqrt{2}\log\frac{c+b}{a})^{2}}\right\},$
so
${\mathrm{\mathbf{P}}}(Z\leq\varepsilon+\operatorname*{{\mathrm{\mathbf{E}}}}Z)\geq 1-\exp\left\{-\frac{n\varepsilon^{2}}{(\sqrt{2}\log\frac{c+b}{a})^{2}}\right\},$
where it follows from $t=n\varepsilon^{2}/(\sqrt{2}\log\frac{c+b}{a})^{2}$ that $\varepsilon=\sqrt{2}\log\left(\frac{c+b}{a}\right)\sqrt{\frac{t}{n}}$. Therefore with probability at least $1-\mathrm{e}^{-t}$,
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert_{\tilde{\mathcal{C}}}\leq\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert_{\tilde{\mathcal{C}}}+\sqrt{2}\log\left(\frac{c+b}{a}\right)\sqrt{\frac{t}{n}}.$
Let $(\varepsilon_{i})_{i\in[n]}$ be i.i.d. Rademacher random variables, independent of $(X_{i})_{i\in[n]}$. By Lemma 22,
$\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert_{\tilde{\mathcal{C}}}\leq 2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}.$
By combining the results above, the following inequality holds with probability at least $1-\mathrm{e}^{-t}$
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert_{\tilde{\mathcal{C}}}\leq 2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}+\sqrt{2}\log\left(\frac{c+b}{a}\right)\sqrt{\frac{t}{n}}.$
Now let $p_{i}=\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-1$, such that $\frac{a}{c+b}\leq p_{i}+1\leq\frac{c+b}{a}$ holds for all $i\in[n]$. Additionally, let $\eta(p)=\log(p+1)$ so that $\eta(0)=0$ and note that for $p\in\left[\frac{a}{c+b}-1,\frac{c+b}{a}-1\right]$, the derivative of $\eta(p)$ is maximal at $p^{*}=\frac{a}{c+b}-1$, and equal to $\eta^{{{}^{\prime}}}(p^{*})=(c+b)/a$. Therefore, $\frac{a}{b+c}\log(p+1)$ is $1$-Lipschitz. By Lemma 20 applied to $\eta(p)$,
$2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$ $=2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\eta(p_{i})\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$
$\leq\frac{4(c+b)}{a}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}-\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$
$\leq\frac{4(c+b)}{a}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}+\frac{4(c+b)}{a}{\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\right\rvert$
$\leq\frac{4(c+b)}{a}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}+\frac{4(c+b)}{a}\frac{1}{\sqrt{n}},$
where the final inequality follows from the following result, proved in Haagerup (1981):
${\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\right\rvert\leq\left({\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\left\{\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\right\}^{2}\right)^{1/2}=\frac{1}{\sqrt{n}}.$
Now, let $\xi_{i}(\tilde{g}_{i})=a\cdot{\tilde{g}(X_{i})}/{\tilde{f}(X_{i})}$, and note that
$\lvert\xi_{i}(u_{i})-\xi_{i}(v_{i})\rvert=\frac{a}{\lvert\tilde{f}(X_{i})\rvert}\lvert u(X_{i})-v(X_{i})\rvert\leq\lvert u(X_{i})-v(X_{i})\rvert.$
By again applying Lemma 20, we have
$\frac{4(c+b)}{a}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$ $\leq\frac{8(c+b)}{a^{2}}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\tilde{g}(X_{i})\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$
$\leq\frac{8(c+b)}{a^{2}}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}g(X_{i})\varepsilon_{i}\right\rVert_{\mathcal{C}}+\frac{8(c+b)}{a^{2}}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}h(X_{i})\varepsilon_{i}\right\rvert$
$\leq\frac{8(c+b)}{a^{2}}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}g(X_{i})\varepsilon_{i}\right\rVert_{\mathcal{C}}+\frac{8(c+b)}{a^{2}}\frac{b}{\sqrt{n}}.$
Applying Lemmas 8 and 23, the following inequality holds for some constant $K>0$:
${\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\sup_{g\in\mathcal{C}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}g(X_{i})\varepsilon_{i}\right\rvert={\operatorname*{{\mathrm{\mathbf{E}}}}}_{\varepsilon}\sup_{\theta\in\Theta}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varphi(X_{i};\theta)\varepsilon_{i}\right\rvert\leq\frac{K}{\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon,$ (13)
and combining the results together, the following inequality holds with probability at least $1-\mathrm{e}^{-t}$:
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert$
$\qquad\leq\frac{8(c+b)K}{a^{2}\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon+\frac{(8b+4a)(c+b)}{a^{2}\sqrt{n}}+\sqrt{2}\log\left(\frac{c+b}{a}\right)\sqrt{\frac{t}{n}},$
$\qquad=\frac{w_{1}}{\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}},$
where $w_{1}$, $w_{2}$, and $w_{3}$ are constants that each depend on some or all of $a$, $b$, and $c$.
Remark 13.
From Lemma 23 we have that $\sigma^{2}_{n}\coloneqq\sup_{f\in\mathscr{F}}P_{n}f^{2}$. To make explicit why $2\sigma_{n}=\left(\sup_{g\in\mathcal{C}}P_{n}g^{2}\right)^{1/2}=2c$, let $\mathscr{F}=\mathcal{C}$ and observe
$\sigma^{2}_{n}=\sup_{g\in\mathcal{C}}P_{n}g^{2}=\sup_{g\in\mathcal{C}}\frac{1}{n}\sum_{i=1}^{n}g(X_{i})^{2}\leq\frac{1}{n}\sum_{i=1}^{n}c^{2}=c^{2}.$
Since our basis functions $\varphi(\cdot,\theta)$ are bounded by $c$, everything greater than $c$ will have value 00 and hence the change from $2c$ to $c$ is inconsequential. However, it can also be motivated by the fact that $\varphi(\cdot,\theta)$ are positive functions.
As highlighted in Remark 12, the full result of Theorem 11 relies on the empirical $L_{2}$ distances $d_{n,x}$ and $d_{n,y}$. In the result of Theorem 9, we make use of the following result to bound $d_{n,x}$ and $d_{n,y}$.
Proposition 14.
By combining Lemmas 16 and 17, the following inequality holds:
$\log N(\mathcal{P},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq\log N_{[]}(\mathcal{P},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq\log N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),$
where $N_{[]}(\mathcal{P},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})$ is the $\varepsilon$-bracketing number of $\mathcal{P}$. Therefore, we have that
$\log N(\mathcal{P},\varepsilon,d_{n,x})\leq\log N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),$
and
$\log N(\mathcal{P},\varepsilon,d_{n,y})\leq\log N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}).$
With this result, we can now prove Theorem 9.
Proof. [of Theorem 9] The notation is the same as in the proof of Theorem 11. The values of the constants may change from line to line.
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)-\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)$
$\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}$
$\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}$
$\quad\qquad+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\quad\qquad-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}$
$\quad\qquad-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}$
$\quad=\left({\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}\right)+\left({\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}\right)$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}\right)$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}\right)$
$\quad\leq 2\sup_{\tilde{g}\in\tilde{\mathcal{C}}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rvert+2\sup_{\tilde{g}\in\tilde{\mathcal{C}}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(Y_{i})}{\tilde{f}(Y_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{g}}{\tilde{f}}\right\rvert$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}\right)$
$\quad\leq 2\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\frac{w^{x}_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon\right\}+\frac{w^{x}_{2}}{\sqrt{n}}+w^{x}_{3}\sqrt{\frac{t}{n}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}$
$\quad\qquad+2\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\frac{w^{y}_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,y})\mathrm{d}\varepsilon\right\}+\frac{w^{y}_{2}}{\sqrt{n}}+w^{y}_{3}\sqrt{\frac{t}{n}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\quad\leq\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}$
$\quad\qquad+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})},$
with probability at least $1-\mathrm{e}^{-t}$, by Theorem 11. Now, we can use (11) from Proposition 6 applied to the target density $f_{k}$, obtaining the following:
$\mathrm{KL}_{h}\left(f_{k}\,||\,f_{k,n}\right)$ $=\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\leq\frac{4a^{-2}c^{2}}{k+2}+\inf_{p\in\mathcal{C}}\mathrm{KL}_{h}\left(f_{k}\,||\,p\right).$
Since by definition we have that $f_{k}\in\mathcal{C}$, $\inf_{p\in\mathcal{C}}\mathrm{KL}_{h}\left(f_{k}\,||\,p\right)=0$, and so with probability at least $1-\mathrm{e}^{-t}$ we have:
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)$ $-\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)$
$\quad\leq\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}+\frac{w_{4}}{k+2}.$ (14)
We can write the overall error as the sum of the approximation and estimation errors as follows. The former is bounded by (10), and the latter is bounded as above in (14). Therefore, with probability at least $1-\mathrm{e}^{-t}$,
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)$ $=[\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)]+[\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)-\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)]$
$\leq\frac{w_{4}}{k+2}+\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}.$
As in Rakhlin et al. (2005), we rewrite the above probabilistic statement as a statement in terms of expectations. To this end, let
$\mathcal{A}\coloneqq\frac{w_{4}}{k+2}+\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}},$
and $\mathcal{Z}\coloneqq\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)$. We have shown ${\mathrm{\mathbf{P}}}\left(\mathcal{Z}\geq\mathcal{A}+w_{3}\sqrt{\frac{t}{n}}\right)\leq\mathrm{e}^{-t}$. Since $\mathcal{Z}\geq 0$,
$\operatorname*{{\mathrm{\mathbf{E}}}}\{\mathcal{Z}\}=\int^{\mathcal{A}}_{0}{\mathrm{\mathbf{P}}}(\mathcal{Z}>s)\mathrm{d}s+\int^{\infty}_{\mathcal{A}}{\mathrm{\mathbf{P}}}(\mathcal{Z}>s)\mathrm{d}s\leq\mathcal{A}+\int_{0}^{\infty}{\mathrm{\mathbf{P}}}(\mathcal{Z}>\mathcal{A}+s)\mathrm{d}s.$
Setting $s=w_{3}\sqrt{\frac{t}{n}}$, we have $t=w_{5}ns^{2}$ and $\operatorname*{{\mathrm{\mathbf{E}}}}\{\mathcal{Z}\}\leq\mathcal{A}+\int_{0}^{\infty}e^{-w_{5}ns^{2}}\mathrm{d}s\leq\mathcal{A}+\frac{w}{\sqrt{n}}$. Hence,
${\operatorname*{{\mathrm{\mathbf{E}}}}}\left\{\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)\right\}-\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)\leq\frac{c_{1}}{k+2}+\frac{c_{2}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N\left(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}\right)\mathrm{d}\varepsilon+\frac{c_{3}}{\sqrt{n}},$
where $c_{1}$, $c_{2}$, and $c_{3}$ are constants that depend on some or all of $a$, $b$, and $c$.
Remark 15.
The approximation error characterises the suitability of the class $\mathcal{C}$, i.e., how well functions in $\mathcal{C}$ are able to estimate a target $f$ which does not necessarily lie in $\mathcal{C}$. The estimation error characterises the error arising from the estimation of the target $f$ on the basis of the finite sample of size $n$.
Proof. [of Corollary 10] Let $\mathcal{X}$ and $\Theta$ be compact and assume the Lipshitz condition given in (12). If $\varphi(x;\cdot)$ is continuously differentiable, then
$\left|\varphi\left(x;\theta\right)-\varphi\left(x;\tau\right)\right|$ $\leq\sum_{k=1}^{d}\left|\frac{\partial\varphi\left(x;\cdot\right)}{\partial\theta_{k}}\left(\theta_{k}^{*}\right)\right|\left|\theta_{k}-\tau_{k}\right|$
$\leq\sup_{\theta^{*}\in\Theta}\left\|\frac{\partial\varphi\left(x;\cdot\right)}{\partial\theta}\left(\theta^{*}\right)\right\|_{1}\left\|\theta-\tau\right\|_{1}.$
Setting
$\varPhi\left(x\right)=\sup_{\theta^{*}\in\Theta}\left\|\frac{\partial\varphi\left(x;\cdot\right)}{\partial\theta}\left(\theta^{*}\right)\right\|_{1},$
we have $\lVert\varPhi\rVert_{\infty}<\infty$. From Lemma 18, we obtain the fact that
$\log N_{[]}\left(\mathcal{P},2\varepsilon\lVert\varPhi\rVert_{\infty},{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}\right)\leq\log N\left(\Theta,\varepsilon,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}\right),$
which by the change of variable $\delta=2\varepsilon\lVert\varPhi\rVert_{\infty}\implies\varepsilon=\delta/2\lVert\varPhi\rVert_{\infty}$ implies
$\log N_{[]}\left(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}\right)\leq\log N\left(\Theta,\frac{\varepsilon}{4\lVert\varPhi\rVert_{\infty}},{\operatorname*{\lVert\,\cdot\,\rVert}}_{1}\right).$
Since $\Theta\subset\mathbb{R}^{d}$, using the fact that a Euclidean set of radius $r$ has covering number
$N\left(r,\varepsilon\right)\leq\left(\frac{3r}{\varepsilon}\right)^{d},$
we have
$\log N\left(\Theta,\frac{\varepsilon}{4\lVert\varPhi\rVert_{\infty}},{\operatorname*{\lVert\,\cdot\,\rVert}}_{1}\right)\leq d\log\left[\frac{12\lVert\varPhi\rVert_{\infty}\mathrm{diam}\left(\Theta\right)}{\varepsilon}\right].$
So
$\int_{0}^{c}\sqrt{\log N\left(\Theta,\frac{\varepsilon}{4\lVert\varPhi\rVert_{\infty}},{\operatorname*{\lVert\,\cdot\,\rVert}}_{1}\right)}\mathrm{d}\varepsilon\leq\int_{0}^{c}\sqrt{d\log\left[\frac{12\lVert\varPhi\rVert_{\infty}\mathrm{diam}\left(\Theta\right)}{\varepsilon}\right]}\mathrm{d}\varepsilon,$
and since $c<\infty$, the integral is finite, as required.
6 Numerical experiments
In this section, we discuss the computability and computation of $\textrm{KL}_{h}$ estimation problems and provide empirical evidence towards the rates obtained in Theorem 9. Namely, we seek to develop a methodology for computing $h$-MLLEs, and to use numerical experiments to demonstrate that the sequence of expected $h$-lifted KL divergences between some density $f$ and a sequence of $k$-component mixture densities from a suitable class $\mathcal{P}$, estimated using $n$ observations does indeed decrease at rates proportional to $1/k$ and $1/\sqrt{n}$, as $k$ and $n$ increase.
6.1 Minorization–Maximization algorithm
One solution for computing problems of kind (4) is to employ an MM algorithm. To do so, we first write the objective of (4) as
$L_{h,n}\left(\psi_{k}\right)=\frac{1}{n}\sum_{i=1}^{n}\left(\log\left\{\sum_{j=1}^{k}\pi_{j}\varphi\left(X_{i};\theta_{j}\right)+h\left(X_{i}\right)\right\}+\log\left\{\sum_{j=1}^{k}\pi_{j}\varphi\left(Y_{i};\theta_{j}\right)+h\left(Y_{i}\right)\right\}\right),$
where $\psi_{k}\in\Psi_{k}=\mathcal{S}_{k}\times\Theta^{k}$. We then require the definition of a minorizer $Q_{n}$ for $L_{h,n}$ on the space $\Psi_{k}$, where $Q_{n}:\Psi_{k}\times\Psi_{k}\rightarrow\mathbb{R}$ is a function with the properties:
  1. (i)
    $Q_{n}\left(\psi_{k},\psi_{k}\right)=L_{h,n}\left(\psi_{k}\right)$, and
  2. (ii)
    $Q_{n}\left(\psi_{k},\chi_{k}\right)\leq L_{h,n}\left(\psi_{k}\right)$,
for each $\psi_{k},\chi_{k}\in\Psi_{k}$. In this context, given a fixed $\chi_{k}$, the minorizer $Q_{n}\left(\cdot,\chi_{k}\right)$ should possess properties that simplify it compared to the original objective $L_{h,n}$. These properties should make the minorizer more tractable and might include features such as parametric separability, differentiability, convexity, among others.
In order to build an appropriate minorizer for $L_{h,n}$, we make use of the so-called Jensen’s inequality minorizer, as detailed in Lange (2016, Sec. 4.3), applied to the logarithm function. This construction results in a minorizer of the form
$Q_{n}\left(\psi_{k},\chi_{k}\right)$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\left\{\tau_{j}\left(X_{i};\chi_{k}\right)\log\pi_{j}+\tau_{j}\left(X_{i};\chi_{k}\right)\log\varphi\left(X_{i};\theta_{j}\right)\right\}$
$+\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\left\{\tau_{j}\left(Y_{i};\chi_{k}\right)\log\pi_{j}+\tau_{j}\left(Y_{i};\chi_{k}\right)\log\varphi\left(Y_{i};\theta_{j}\right)\right\}$
$+\frac{1}{n}\sum_{i=1}^{n}\left\{\gamma\left(X_{i};\chi_{k}\right)\log h\left(X_{i}\right)+\gamma\left(Y_{i};\chi_{k}\right)\log h\left(Y_{i}\right)\right\}$
$-\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\left\{\tau_{j}\left(X_{i};\chi_{k}\right)\log\tau_{j}\left(X_{i};\chi_{k}\right)+\tau_{j}\left(Y_{i};\chi_{k}\right)\log\tau_{j}\left(Y_{i};\chi_{k}\right)\right\}$
$-\frac{1}{n}\sum_{i=1}^{n}\left\{\gamma\left(X_{i};\chi_{k}\right)\log\gamma\left(X_{i};\chi_{k}\right)+\gamma\left(Y_{i};\chi_{k}\right)\log\gamma\left(Y_{i};\chi_{k}\right)\right\}$
where
$\gamma\left(X_{i};\psi_{k}\right)=h\left(X_{i}\right)/\left\{\sum_{j=1}^{k}\pi_{j}\varphi\left(X_{i};\theta_{j}\right)+h\left(X_{i}\right)\right\}$
and
$\tau_{j}\left(X_{i};\psi_{k}\right)=\pi_{j}\varphi\left(X_{i};\theta_{j}\right)/\left\{\sum_{j=1}^{k}\pi_{j}\varphi\left(X_{i};\theta_{j}\right)+h\left(X_{i}\right)\right\}.$
Observe that $Q_{n}\left(\cdot,\chi_{k}\right)$ now takes the form of a sum-of-logarithms, as opposed to the more challenging log-of-sum form of $L_{h,n}$. This change produces a functional separation of the elements of $\psi_{k}$.
Using $Q_{n}$, we then define the MM algorithm via the parameter sequence $\left(\psi_{k}^{\left(s\right)}\right)_{s\in\mathbb{N}}$, where
$\psi_{k}^{\left(s\right)}=\underset{\psi_{k}\in\Psi_{k}}{\operatorname*{arg\,max\,}}~{}Q_{n}\left(\psi_{k},\psi_{k}^{\left(s-1\right)}\right),$ (15)
for each $s>0$, and where $\psi_{k}^{\left(0\right)}$ is user chosen and is typically referred to as the initialization of the algorithm. Notice that for each $s$, (15) is a simpler optimization problem than (4). Writing $\psi_{k}^{\left(s\right)}=\left(\pi_{1}^{\left(s\right)},\dots,\pi_{k}^{\left(s\right)},\theta_{1}^{\left(s\right)},\dots,\theta_{k}^{\left(s\right)}\right)$, we observe that (15) simplifies to the separated expressions:
$\pi_{j}^{\left(s\right)}=\frac{\sum_{i=1}^{n}\left\{\tau_{j}\left(X_{i};\psi_{k}^{\left(s-1\right)}\right)+\tau_{j}\left(Y_{i};\psi_{k}^{\left(s-1\right)}\right)\right\}}{\sum_{i=1}^{n}\sum_{l=1}^{k}\left\{\tau_{l}\left(X_{i};\psi_{k}^{\left(s-1\right)}\right)+\tau_{l}\left(Y_{i};\psi_{k}^{\left(s-1\right)}\right)\right\}}$
and
$\theta_{j}^{\left(s\right)}=\underset{\theta_{j}\in\Theta}{\operatorname*{arg\,max\,}}\frac{1}{n}\sum_{i=1}^{n}\left\{\tau_{j}\left(X_{i};\psi_{k}^{\left(s-1\right)}\right)\log\varphi\left(X_{i};\theta_{j}\right)+\tau_{j}\left(Y_{i};\psi_{k}^{\left(s-1\right)}\right)\log\varphi\left(Y_{i};\theta_{j}\right)\right\},$
for each $j\in\left[k\right]$.
A noteworthy property of the MM sequence $\left(\psi_{k}^{\left(s\right)}\right)_{s\in\mathbb{N}}$ is that it generates an increasing sequence of objective values, due to the chain of inequalities
$L_{h,n}\left(\psi_{k}^{\left(s-1\right)}\right)=Q_{n}\left(\psi_{k}^{\left(s-1\right)},\psi_{k}^{\left(s-1\right)}\right)\leq Q_{n}\left(\psi_{k}^{\left(s\right)},\psi_{k}^{\left(s-1\right)}\right)\leq L_{h,n}\left(\psi_{k}^{\left(s\right)}\right),$
where the equality is due to property (i) of $Q_{n}$, the first in equality is due to the definition of $\psi_{k}^{\left(s\right)}$, and the second inequality is due to property (ii) of $Q_{n}$. This provides a kind of stability and regularity to the sequence $\left(L_{h,n}\left(\psi_{k}^{\left(s\right)}\right)\right)_{s\in\mathbb{N}}$.
Of course, we can provide stronger guarantees under additional assumptions. Namely, assume that (iii) $\Psi_{k}\subset\varPsi_{k}$, where $\varPsi_{k}$ is an open set in a finite dimensional Euclidean space on which $L_{h,n}$ and $Q_{n}\left(\cdot,\chi_{k}\right)$ is differentiable, for each $\chi_{k}\in\Psi_{k}$. Then, under assumptions (i)–(iii) regarding $L_{h,n}$ and $Q_{n}$, and due to the compactness of $\Psi_{k}$ and the continuity of $Q_{n}$ on $\Psi_{k}\times\Psi_{k}$, Razaviyayn et al. (2013, Cor. 1) implies that $\left(\psi_{k}^{\left(s\right)}\right)_{s\in\mathbb{N}}$ converges to the set of stationary points of $L_{h,n}$ in the sense that
$\lim_{s\rightarrow\infty}\inf_{\psi_{k}^{*}\in\Psi_{k}^{*}}\left\|\psi_{k}^{\left(s\right)}-\psi_{k}^{*}\right\|_{2}=0$
where,
$\Psi_{k}^{*}=\left\{\psi_{k}^{*}\in\Psi_{k}:\left.\frac{\partial L_{h,n}}{\partial\psi_{k}}\right|_{\psi_{k}=\psi_{k}^{*}}=0\right\}.$
More concisely, we say that the sequence $\left(\psi_{k}^{\left(s\right)}\right)_{s\in\mathbb{N}}$ globally converges to the set of stationary points $\Psi_{k}^{*}$.
6.2 Experimental setup
Towards the task of demonstrating empirical evidence of the rates in Theorem 9, we consider the family of beta distributions on the unit interval $\mathcal{X}=\left[0,1\right]$ as our base class (i.e., (6)) to estimate a pair of target densities
$f_{1}\left(x\right)=\frac{1}{2}\chi_{\left[0,2/5\right]}\left(x\right)+\frac{1}{2}\chi_{\left[3/5,1\right]}\left(x\right),$
and
$f_{2}\left(x\right)=\chi_{\left[0,1\right]}\left(x\right)\begin{cases}2-4x&\text{if }x\leq 1/2,\\ -2+4x&\text{if }x>1/2,\end{cases}$
where $\chi_{\mathcal{A}}$ is the characteristic function that takes value $1$ if $x\in\mathcal{A}$ and 00, otherwise. Note that neither $f_{1}$ nor $f_{2}$ are in $\mathcal{C}$. In particular, $f_{1}\left(x\right)=0$ when $x\in\left(\nicefrac{{2}}{{5}},\nicefrac{{3}}{{5}}\right)$, and $f_{2}(x)=0$ when $x=1/2$, and hence neither densities are bounded away from 0, on $\mathcal{X}$. Thus, the theory of Rakhlin et al. (2005) cannot be applied to provide bounds for the expected KL divergence between MLEs of beta mixtures and the pair of targets. We visualize $f_{1}$ and $f_{2}$ in Figure 1.
Figure 1: Simulation target densities $f_{1}$ (solid line) and $f_{2}$ (dashed line).
To observe the rate of decrease of the $h$-lifted KL divergence between the targets and respective sequences of $h$-MLLEs, we conduct two experiments E1 and E2. In E1, our target density is set to $f_{1}$ and $h_{1}=\beta\left(\cdot;1/2,1/2\right)$. For each $n\in\left\{2^{10},\dots,2^{15}\right\}$ and $k\in\left\{2,\dots,8\right\}$, we independently simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$ with each $X_{i}$ and $Y_{i}$ ($i\in\left[n\right]$), i.i.d., from the distributions characterized by $f_{1}$ and $h_{1}$, respectively. In E2, we target $f_{2}$ with $h$-MLLEs over the same ranges of $k$ and $n$, but with $h_{2}=\beta\left(\cdot;1,1\right)$–the density of the uniform distribution. For each $k$ and $n$, we simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$ respectively from distributions characterized by $f_{2}$ and $h_{2}$.
In both experiments, we simulate $r=50$ replicates of each $\left(k,n\right)$-scenario and compute the corresponding $h$-MLLEs, $\left(f_{k,n,l}\right)_{l\in[r]}$, using the previously described MM algorithm. For each $l\in\left[r\right]$, we compute the corresponding negative log $h$-lifted likelihood between the target $f$ and $f_{k,n,l}$:
$K_{k,n,l}=-\int_{\mathcal{X}}\left(f+h\right)\log\left(f_{k,n,l}+h\right)\mathrm{d}\mu$
to assess the rates, and note that
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n,l}\right)=\int_{\mathcal{X}}\left(f+h\right)\log\left(f+h\right)\mathrm{d}\mu+K_{k,n,l},$
where the prior term is a constant with respect to $k$ and $n$.
To analyze the sample of $7\times 6\times 50=2100$ observations of relationship between the values $\left(K_{k,n,l}\right)_{l\in\left[r\right]}$ and the corresponding values of $k$ and $n$, we use non-linear least squares (Amemiya, 1985, Sec. 4.3) to fit the regression relationship:
$\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]=a_{0}+\frac{a_{1}}{\left(k+2\right)^{b_{1}}}+\frac{a_{2}}{n^{b_{2}}}.$ (16)
We obtain $95\%$ asymptotic confidence intervals for the estimates of the regression parameters $a_{0},a_{1},a_{2},b_{1},b_{2}\in\mathbb{R}$, under the assumption of potential mis-specification of (16), by using the sandwich estimator for the asymptotic covariance matrix (cf. White 1982). We include the code for these experiments in Appendix B.
6.3 Results
We report the estimates along with $95\%$ asymptotic confidence intervals for the parameters of (16) for E1 and E2 in Table 1. Plots of the average negative log $h$-lifted likelihood values by sample sizes $n$ and numbers of components $k$ are provided in Figure 2.
Table 1: Estimates of parameters for fitted relationships (with $95\%$ confidence intervals) between negative log $h$-lifted likelihood values, sample size and number of mixture components for experiments E1 and E2.
E1 $a_{0}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$
Est. $-1.68$ $0.73$ $6.80$ $1.87$ $0.99$
95% CI $\left(-1.68,-1.67\right)$ $\left(0.68,0.78\right)$ $\left(1.24,12.36\right)$ $\left(1.81,1.93\right)$ $\left(0.87,1.11\right)$
E2 $a_{0}$ $a_{1}$ $a_{2}$ $b_{1}$ $b_{2}$
Est. $-1.47$ $1.49$ $6.75$ $4.36$ $1.07$
95% CI $\left(-1.48,-1.47\right)$ $\left(0.58,2.41\right)$ $\left(2.17,11.32\right)$ $\left(3.91,4.81\right)$ $\left(0.97,1.16\right)$
Figure 2: Average negative log $h$-lifted likelihood values by sample sizes $n$ and numbers of components $k$ for experiments E1 and E2.
From Table 1, we observe that $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ decreases with both $n$ and $k$ in both simulations, and that the rates at which the averages decrease are faster than anticipated by Theorem 9, with respect to both $n$ and $k$. We can visually confirm the decreases in the estimate of $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ via Figure 2. In both E1 and E2, the rate of decrease over the assessed range of $n$ is approximately proportional to $1/n$, as opposed to the anticipated rate of $1/\sqrt{n}$, whereas the rate of decrease in $k$ is far larger, at approximately $1/k^{1.87}$ for E1 and $1/k^{4.36}$ for E2.
These observations provide empirical evidence towards the fact that the rate of decrease of $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ is at least $1/k$ and $1/\sqrt{n}$, respectively, for $k$ and $n$, at least over the simulation scenarios. These fast rates of fit over small values of $n$ and $k$ may be indicative of a diminishing returns of fit phenomenon, as discussed in Cadez and Smyth (2000) or the so-called elbow phenomenon (see, e.g., Ritter 2014, Sec. 4.2), whereupon the rate of decrease in average loss for small values of $k$ is fast and becomes slower as $k$ increases, converging to some asymptotic rate. This is also the reason why we do not include the outcomes when $k=1$, as the drop in $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ between $k=1$ and $k=2$ is so dramatic that it makes our simulated data ill-fitted by any model of form (16). As such, we do not view Theorem 9 as being pessimistic in light of these phenomena, as it applies uniformly over all values of $k$ and $n$.
7 Conclusion
The estimation of probability densities using finite mixtures from some base class $\mathcal{P}$ appears often in machine learning and statistical inference as a natural method for modeling underlying data generating processes. In this work, we sought to provide novel generalization bounds for such mixture estimators. To this end, we introduce the family of $h$-lifted KL divergences for densities on compact supports, within the family of Bregman divergences, which correspond to risk functions that can be bounded, even when densities in the class $\mathcal{P}$ are not bounded away from zero, unlike the standard KL divergence. Unlike the least-squares loss, the corresponding maximum $h$-likelihood estimation problem can be computed via an MM algorithm, mirroring the availability of EM algorithms for the maximum likelihood problem corresponding to the KL divergence. Along with our derivations of generalization bounds that achieve the same rates as the best-known bounds for the KL divergence and least square loss, we also provide numerical evidence towards the correctness of these bounds in the case when $\mathcal{P}$ corresponds to beta distribution densities.
Aside from beta distributions, mixture densities on compact supports that can be analysed under our framework appear frequently in the literature. For supports on compact Euclidean subset, examples include mixtures of Dirichlet distributions (Fan et al., 2012) and bivariate binomial distributions (Papageorgiou and David, 1994). Alternatively, one can consider distributions on compact Euclidean manifolds, such as mixtures of Kent (Peel et al., 2001) distributions and von Mises–Fisher distributions (Banerjee et al., 2005, Ng and Kwong, 2022). We defer investigating the practical performance of the maximum $h$-lifted likelihood estimators and accompanying theory for such models to future work.
Appendix A The $h$-lifted KL divergence as a Bregman divergence
Let $\tilde{u}=u+h$, so that $\phi(u)=\tilde{u}\log(\tilde{u})-\tilde{u}+1$. Then $\phi^{\prime}(u)=\log(\tilde{u})$, and
$D_{\phi}(p\,||\,q)$ $=\int_{X}\{\tilde{p}\log(\tilde{p})-\tilde{p}-1\}-\{\tilde{q}\log(\tilde{q})-\tilde{q}-1\}-\log(\tilde{q})(p-q)\mathrm{d}\mu$
$=\int_{\mathcal{X}}\tilde{p}\log(\tilde{p})-\tilde{q}\log(\tilde{q})-p\log(\tilde{q})+q\log(\tilde{q})\mathrm{d}\mu$
$=\int_{\mathcal{X}}\{p+h\}\log(\tilde{p})-\{q+h\}\log(\tilde{q})-p\log(\tilde{q})+q\log(\tilde{q})\mathrm{d}\mu$
$=\int_{\mathcal{X}}\{p+h\}\log{\frac{p+h}{q+h}}\mathrm{d}\mu=\mathrm{KL}_{h}\left(p\,||\,q\right).$
Appendix B Experimental results
The code for all simulations and analyses of Experiment 1 and 2 is available in both the R and Python programming languages. The experiments and analyses presented were conducted in R. The figures were created in Python. The code repository is available here: https://github.com/hiendn/LiftedLikelihood.
Appendix C Technical results
Here we collect some technical results that are required in our proofs but appear elsewhere in the literature. In some places, notation may be modified from the original text to keep with the established conventions herein.
Lemma 16 (Kosorok, 2007. Lem 9.18).
[…]
$N(\mathscr{F},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq N_{[]}(\mathscr{F},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})$
for all $\varepsilon>0$.
Lemma 17 (Kosorok, 2007. Lem 9.22).
[…]
$\log N_{[]}(\mathscr{F},2\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})\leq\log N(\mathscr{F},\varepsilon,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),$
for all $\varepsilon>0$.
Lemma 18 (Kosorok, 2007. Thm 9.23).
Let =$\mathscr{F}=\{f_{t}:t\in T\}$ be a function class satisfying
$\lvert f_{s}(x)-f_{t}(x)\rvert\leq d(s,t)F(x),$
[…]
$N_{[]}(\mathscr{F},2\varepsilon\lVert F\rVert,\operatorname*{\lVert\,\cdot\,\rVert})\leq N(T,\varepsilon,d).$
Lemma 19 (Shalev-Shwartz and Ben-David (2014), Lem 26.7).
Let $A$ be a subset of $\mathbb{R}^{m}$ and let
$A^{\prime}=\left\{\sum_{j=1}^{n}\alpha_{j}\mathbf{a}_{j}\mid n\in\operatorname*{\mathbb{N}},\mathbf{a}_{j}\in A,\alpha_{j}\geq 0,\lVert\mathbf{\alpha}\rVert_{1}=1\right\}.$
Then, $\mathcal{R}_{n}(A^{\prime})=\mathcal{R}_{n}(A)$, i.e., both $A$ and $A^{\prime}$ have the same Rademacher complexity.
Lemma 20 (van de Geer, 2016, Thm. 16.2).
Let $(X_{i})_{i\in[n]}$ be non-random elements of $\mathcal{X}$ and let $\mathscr{F}$ be a class of real-valued functions on $\mathcal{X}$. If $\varphi_{i}:\operatorname*{\mathbb{R}}\to\operatorname*{\mathbb{R}}$, $i\in[n]$, are functions vanishing at zero that satisfy for all $u,v\in\operatorname*{\mathbb{R}}$,
$\lvert\varphi_{i}(u)-\varphi_{i}(v)\rvert\leq\lvert u-v\rvert,$
then we have
${\mathrm{\mathbf{E}}}\left\{\left\lVert\sum_{i=1}^{n}\varphi_{i}(f(X_{i}))\varepsilon_{i}\right\rVert_{\mathscr{F}}\right\}\leq 2{\mathrm{\mathbf{E}}}\left\{\left\lVert\sum_{i=1}^{n}f(X_{i})\varepsilon_{i}\right\rVert_{\mathscr{F}}\right\}.$
Lemma 21 (McDiarmid, 1998, Thm. 3.1 or McDiarmid, 1989).
Suppose $(X_{i})_{i\in[n]}$ are independent random variables and let $Z=g(X_{1},\ldots,X_{n})$, for some function $g$. If $g$ satisfies the bounded difference condition, that is there exists constant $c_{j}$ such that for all $j\in[n]$ and all $x_{1},\ldots,x_{j},x_{j}^{\prime},\ldots,x_{n}$,
$\lvert g(x_{1},\ldots,x_{j-1},x_{j},x_{j+1},\ldots,x_{n})-g(x_{1},\ldots,x_{j-1},x_{j}^{\prime},x_{j+1},\ldots,x_{n})\rvert\leq c_{j},$
then
${\mathrm{\mathbf{P}}}(Z-{\mathrm{\mathbf{E}}}Z\geq t)\leq\exp\left\{\frac{-2t^{2}}{\sum_{j=1}^{n}c_{j}^{2}}\right\}.$
Lemma 22 (van der Vaart and Wellner, 1996, Lem. 2.3.1).
Let $\mathfrak{R}(f)=\mathrm{\mathbf{E}}f$ and $\mathfrak{R}_{n}(f)=n^{-1}\sum_{i=1}^{n}f(X_{i})$. If $\varPhi:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ is a convex function, then the following inequality holds for any class of measurable functions $\mathscr{F}$:
${\mathrm{\mathbf{E}}}\varPhi\left(\left\lVert\mathfrak{R}(f)-\mathfrak{R}_{n}(f)\right\rVert_{\mathscr{F}}\right)\leq{\mathrm{\mathbf{E}}}\varPhi\left(2\left\lVert R_{n}(f)\right\rVert_{\mathscr{F}}\right),$
where $R_{n}(f)$ is the Rademacher process indexed by $\mathscr{F}$. In particular, since the identity map is convex,
${\mathrm{\mathbf{E}}}\left\{\lVert\mathfrak{R}(f)-\mathfrak{R}_{n}(f)\rVert_{\mathscr{F}}\right\}\leq 2{\mathrm{\mathbf{E}}}\left\{\lVert R_{n}(f)\rVert_{\mathscr{F}}\right\}.$
Lemma 23 (Koltchinskii, 2011, Thm. 3.11).
Let $d_{n}$ be the empirical distance
$d_{n}^{2}(f_{1},f_{2})=\frac{1}{n}\sum_{i=1}^{n}(f_{1}(X_{i})-f_{2}(X_{i}))^{2}$
and denote by $N(\mathscr{F},\varepsilon,d_{n})$ the $\varepsilon$-covering number of $\mathscr{F}$. Let $\sigma_{n}^{2}\coloneqq\sup_{f\in\mathscr{F}}P_{n}f^{2}$. Then the following inequality holds
${\mathrm{\mathbf{E}}}\left\{\lVert R_{n}(f)\rVert_{\mathscr{F}}\right\}\leq\frac{K}{\sqrt{n}}\mathrm{\mathbf{E}}\int_{0}^{2\sigma_{n}}\log^{1/2}N(\mathscr{F},\varepsilon,d_{n})\mathrm{d}\varepsilon$
for some constant $K>0$.